OpenStudy (anonymous):

(3a)/(a^2+a-12) - (4+a)/(9-a^2)

OpenStudy (anonymous):

\[\frac{ 3a }{ a^{2} + a -12 }-\frac{ 4+a }{ 9-a^2 }\] so what you need to do first is find a common denominator

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

skull can you help me?

OpenStudy (skullpatrol):

Find the product of the greatest power of each factor occurring in the denominators.

OpenStudy (anonymous):

What do you mean @skullpatrol

OpenStudy (ybarrap):

divide a^2 + a -12 by one of the factors of 9-a^2. 9-a^2=(3+a)(3-a) Let's try a-3, we'll add the negative back in later dividing a^2 + a -12 by a-3, we get a+4 so a-3, a+3, and a+4 are common factors, this is our common denominator So solution is \[(16 + a (17 + 4 a))/((-3 + a) (3 + a) (4 + a))\]

OpenStudy (anonymous):

Ohhh. Thank you!